In manufacturing a medical equipment, a firm incurs costs of two types. Fixed annual costs of $ 145,500 are incurred regardless of the number of units produced. In addition, each unit produced costs the firm $7. If C equals total annual cost in dollars and x equals the number of units produced during a year; a) Develop the linear function C = f(x) which expresses annual cost. b) Calculate f (320)? What does f (320) represent? c) What is the y-intercept in the linear function that you have formed in part a and what does it signify?


1
Expert's answer
2021-04-15T07:27:01-0400

Answer: "a) C(x)=145500+7x;"

"b) C(320000)=145500+7\sdot320000=2385500"

It is needed $2,385,500 to produce 320,000 units.

"c)b=145500 -" y-intercept.

It is fixed annual costs are incurred regardless of the number of units produced.